# An Accurate Efficient Analytic Model of Fidelity under Depolarizing Noise oriented to Large Scale Quantum System Design

### v1

@@ 4.1 Fidelity Estimation Algorithm
Building upon the theoretical findings about fidelity loss under depolarizing noise presented in Section 2.1, we present the fidelity estimation algorithm used in the previous results. This algorithm predicts the expected fidelity of a quantum circuit based on its structure and the calibration data of the quantum processor on which the circuit is being executed, under the assumption that each gate introduces depolarizing noise. The algorithm is summarized in Algorithm 1.
@@
Leveraging the unitary invariance of both the depolarizing channel and quantum fidelity, all gates in the circuit are effectively replaced with fidelity-equivalent depolarizing channels. The proposed algorithm initializes each qubit with a fidelity of 1 and iteratively reduces the fidelity based on the depolarizing channels applied to each quantum state. The global fidelity is determined using the multiplicativity under tensor product property of fidelity. Specifically, the circuit’s overall fidelity ($F_{circ}$) is obtained as the product of the individual qubit-wise fidelities ($F_{q}$):
@@
The results in Figure 3 reveal that the fidelity range determined by $p_{ent}=0$ and $p_{ent}=1$ is relatively narrow, in contrast to QVA, where the hyperparameter $w$ can cause fidelity estimates to vary by as much as 0.8. This demonstrates that an intermediate value, such as $p_{ent}=0.5$, can reasonably estimate the fidelity. Consequently, the inability to precisely determine the level of entanglement does not significantly impact the accuracy of the fidelity estimation.
To validate the approach presented in this work, it is essential to incorporate the effects of coherence errors [20, 3, 4] into the fidelity estimation model, as these errors are a significant source of decoherence in current quantum hardware. The exact decoherence induced by relaxation ($T_{1}$) and dephasing ($T_{2}$) depends on the particular quantum state (i.e., they are not unitary-invariant), which can only be fully assessed by simulating the quantum state. Instead, we assume a worst-case scenario where qubits undergo maximal decoherence, thus providing an upper bound on fidelity loss.
The resulting upper bound on fidelity loss is incorporated into Algorithm 1 by scaling the qubit-wise fidelity after the execution of each layer (i.e., a set of gates that can be applied simultaneously on independent qubits) of the circuit to account for coherence errors. Specifically, the relaxation time $T_{1}$ and dephasing time $T_{2}$ for each qubit are obtained from the device calibration data, and the execution time for each layer is estimated based on the duration of the gates that can be executed in parallel. After each layer, the fidelity of each qubit is updated as:
@@ 4.2 Quantum Processor Execution
This section details the experimental setup and workflow for evaluating circuit success rates and comparing fidelity estimation methods. Figure 7 provides an overview of the entire workflow.
Figure 7: Workflow for obtaining fidelity predictions from the four estimation methods (Qiskit, ESP, QVA, proposed model, and proposed $+T_{1,2}$) and for determining the success rate from real-device executions. The process includes circuit selection, compilation to match device-specific constraints, fidelity estimation based on noise models and calibration data, execution on a real quantum processor, and measurement error mitigation to compute the success rate. This workflow ensures a consistent comparison of predictions with hardware performance.
@@
We selected circuits ranging from 2 to 8 qubits ($q$) sourced from MQT Bench [49]. Each circuit $\mathcal{C}$ was concatenated with its inverse $\mathcal{C}^{-1}$, hence ensuring the final quantum state being equal to the initial state, the all-zero state $|0\rangle^{\otimes q}$ in our case, with a corresponding ideal measurement outcome of ’0’×$\times q$.
@@
Before running the circuits on the actual quantum device, they went through the compilation process, which decomposed gates into native gates supported by the device [50, 51], mapped logical qubits to physical qubits [52], and added routing operations [53, 54] to meet hardware connectivity constraints [55]. We used Qiskit [36] for all stages of this compilation process, ensuring consistency across all the evaluated circuits. The result is a compiled circuit optimized for execution on the chosen quantum processor.
@@
For each estimation method, fidelity predictions were made using the compiled circuit and calibration data from the quantum processor. For Qiskit simulations, noise models were directly obtained from the Qiskit library [36], tailored to the target device, and simulations were conducted using the density_matrix method to account for mixed-state evolution under noise.
@@
Experiments were conducted on an IBM Eagle r3 processor [37] with 127 qubits. The compiler selected the most reliable physical qubits for each circuit. After execution, qubits were measured, and readout errors were mitigated using calibration data and assuming uncorrelated measurement errors between the qubits. The success rate is calculated as the ratio of correctly measured outcomes (’0’×$\times q$) over the total number of executions (2048 per circuit).
@@
This setup establishes a robust framework for evaluating the performance of the real quantum device against fidelity predictions from various estimation methods, providing valuable insights into the accuracy and dependability of each model.
5 Data Availability
The datasets generated and/or analysed during the current study are available in the GitHub repository, https://github.com/escofetpau/Analytic-Model-of-Fidelity-under-Depolarizing-Noise.
6 Code Availability
@@
The underlying code for this study is available in GitHub and can be accessed via this link https://github.com/escofetpau/Analytic-Model-of-Fidelity-under-Depolarizing-Noise.
7 Acknowledgements
@@ Appendix A Robustness of Fidelity Estimation Across Multiple Hardware Executions
Figure 8: Comparison of the fidelity prediction accuracy for five different estimation models, each incorporating coherence errors ($T_{1}$ and $T_{2}$). The figure consists of five rows, one per model (including QVA with $w=0$), and six columns, each corresponding to an independent experiment conducted on different days using IBM Q Eagle processors (IBM Q Kyiv, IBM Q Sherbrooke, and IBM Q Brisbane). Each cell represents the fidelity error for a given model in a specific experiment. The $R^{2}$ metric is reported for each model, quantifying the correlation between predicted and measured fidelities. Across all experiments, the proposed model consistently achieves the highest $R^{2}$ values, ranging from $0.87$ to $0.94$, demonstrating superior predictive accuracy and robustness.

### v2

@@ 4.1 Fidelity Estimation Algorithm
We propose the fidelity estimation algorithm based on the theoretical findings about fidelity loss under depolarizing noise presented in Section 3. This algorithm predicts the expected fidelity of a quantum circuit based on its structure and the calibration data of the quantum processor on which the circuit is being executed, assuming that each gate introduces depolarizing noise. It is summarized in Algorithm 1.
@@
By leveraging the unitary invariance of both the depolarizing channel and quantum fidelity, all gates in the circuit are effectively replaced with fidelity-equivalent depolarizing channels. Each qubit’s independent fidelity is initialized to 1 and iteratively reduced based on the depolarizing channels applied to each quantum state. The global fidelity is determined by the multiplicativity under tensor product property of fidelity. Specifically, the circuit’s overall fidelity ($F_{circ}$) is obtained as the product of the individual qubit-wise fidelities ($F_{q}$):
@@
It is essential to incorporate the effects of coherence errors [26, 3, 4] into the fidelity estimation model to validate the approach presented in this work against real quantum hardware, as these errors are a significant source of decoherence in current quantum hardware. The exact decoherence induced by relaxation ($T_{1}$) and dephasing ($T_{2}$) depends on the particular quantum state (i.e., they are not unitary-invariant), which can only be fully assessed by simulating the whole quantum circuit. Instead, we assume a worst-case scenario where qubits undergo maximal decoherence, thus providing an upper bound on fidelity loss.
The resulting decoherence upper bound is incorporated into Algorithm 1 by scaling the qubit-wise fidelity after the execution of each layer of the circuit (i.e., a set of gates that can be applied simultaneously on independent qubits) to account for coherence errors. Specifically, the relaxation time $T_{1}$ and dephasing time $T_{2}$ for each qubit are obtained from the device calibration data, and the execution time for each layer is estimated based on the duration of the gates that can be executed in parallel. After each layer, the fidelity of each qubit (even the ones that remained idling during that layer) is scaled as:
@@ 4.2 Quantum Processor Execution
This subsection details the experimental setup and workflow for evaluating circuit success rates and comparing fidelity estimation methods. Figure 3 provides an overview of the entire workflow.
@@
We selected quantum circuits ranging from 2 to 8 qubits, obtained from the MQT Bench [45]. Each circuit $\mathcal{C}$ was concatenated with its inverse $\mathcal{C}^{-1}$, ensuring that the final quantum state is equal to the initial one, the all-zero state $|0\rangle^{\otimes q}$ in our case, with a corresponding ideal measurement outcome of ’0’×$\times q$.
@@
Before running the circuits on the actual quantum device, they went through the compilation process, which decomposed gates into native gates supported by the device [46, 47], mapped logical qubits to physical qubits [48], and added routing operations [49, 50] to meet hardware connectivity constraints [51]. We used Qiskit [36] for all stages of the compilation process, ensuring consistency across all the evaluated circuits. The result is a compiled circuit optimized for execution on the chosen quantum processor.
Figure 3: Workflow for obtaining fidelity predictions from the four estimation methods (Qiskit, ESP, QVA, proposed model, and proposed $+T_{1,2}$) and determining the success rate from real-device executions. The process includes circuit selection, compilation to match device-specific constraints, fidelity estimation based on noise models and calibration data, execution on a real quantum processor, and measurement error mitigation to compute the success rate. This workflow ensures a consistent comparison of predictions with hardware performance.
@@
For each estimation method, fidelity predictions were made using the compiled circuit and calibration data from the quantum processor at the execution time. For Qiskit simulations, noise models were directly obtained from the Qiskit library [36], tailored to the target device, and simulations were conducted using the density_matrix method to account for mixed-state evolution under noise.
@@
Experiments were conducted on an IBM Eagle r3 processor [37] with 127 qubits. The compiler selected the most reliable subset of physical qubits for each circuit. After execution, qubits were measured, and readout errors were mitigated using calibration data and assuming uncorrelated measurement errors between the qubits. The success rate is calculated as the ratio of correctly measured outcomes (’0’×$\times q$) over the total number of executions (2048 per circuit).
@@
This setup establishes a robust framework for evaluating the performance of the real quantum device against fidelity predictions from various estimation methods, providing valuable insights into each model’s accuracy and dependability.
5 Results
Having validated the model through simulation in Section 3.1, we now assess its performance against real quantum hardware.
The fidelity for each compiled circuit in the benchmark is estimated using the proposed model alongside other methods, including Qiskit simulation (using noise models associated with the used processor), ESP, and QVA. We compare these predictions to the success rates obtained from executions on quantum hardware after readout error mitigation. Additionally, a version of each model incorporating coherence errors ($T_{1}$ and $T_{2}$), as detailed in Section 4.1, is also considered to provide a more comprehensive assessment of actual executions. Since quantum hardware is subject to additional sources of error beyond those captured by the considered models, achieving an accurate prediction by considering only gate errors is impossible, making the inclusion of $T_{1}$ and $T_{2}$ essential for a robust validation against quantum hardware.
Figure 4 illustrates the measured success rate in the quantum device (IBM Q Eagle processor) after readout error mitigation (on the $y$-axis) versus the predicted fidelity (on the $x$-axis) for each estimation method. In this plot, a perfect prediction would place the data point along the diagonal. For the QVA and proposed models, horizontal lines indicate the influence of the hyperparameters in each algorithm ($w$ for QVA and $p_{ent}$ for the proposed model), with the reported point representing the average value between the extremes of these parameters. Most gate-error models tend to estimate a higher fidelity than the actual observation on the real device, which is expected since the models do not account for all types of errors present in actual hardware (e.g., coherence or crosstalk errors). This effect is evident when incorporating coherence errors into the estimation models, as it demonstrates a notable improvement in aligning predictions with the observed success rates for all models except for QVA, where it leads to a worse prediction, estimating a lower fidelity than the observed one.
Figure 4: Comparison of predicted versus measured fidelities for all estimation models. Each scatter plot illustrates the predicted fidelity ($x$-axis) against the measured success rate ($y$-axis), with the diagonal line ($x=y$) indicating perfect predictions. For the QVA and the proposed model, horizontal lines indicate the range of predicted fidelities arising from the variability of their hyperparameters $w$ and $p_{ent}$, respectively.
Interestingly, QVA sometimes predicts a lower fidelity than the observed success rate despite not including coherence or crosstalk errors in its model. This discrepancy arises because QVA decreases the global fidelity twice for each two-qubit gate, doubling the reduction in fidelity compared to the actual gate error. When coherence errors are incorporated into the QVA model, the predicted fidelities decrease even further, as shown in the lower row of Figure 4. Moreover, as mentioned before, the long horizontal lines for the QVA in Figure 4 highlight the significant influence that the hyperparameter for the cross-error ($w$) has on its predictions.
Across all estimation techniques, we observe that, when only taking into account gate errors (upper row), for circuits with either very few or very many gates, the predictions tend to be relatively accurate. The explanation lies in the simplicity of estimating fidelity in these cases. For circuits with a very low number of gates, the fidelity is typically close to one, and fewer interactions among qubits reduce cross-talk errors, which are not considered by any of the evaluated models. Additionally, coherence noise effects are less significant due to the shorter execution times of such circuits. Conversely, the fidelity generally approaches zero for circuits with a high number of gates, causing the absolute difference between the predicted and real success rates to decrease.
To assess the accuracy of the proposed prediction models, we compute several regression metrics typically used to evaluate the performance of predictive models. These include the Mean Absolute Error (MAE) [52], which measures the average absolute difference between predicted and actual values; the Mean Squared Error (MSE) [53], which measures the average of the squares of the errors, that is, the average squared difference between the estimated values and the actual value; and the $R^{2}$ Score [54], which quantifies the proportion of variance in the observed data that is predictable from the model. Finally, the Pearson Correlation [55] coefficient is used to assess the linear relationship between the predicted and actual values, where a higher value indicates a stronger linear relationship. Additionally, the results with the cross-error parameter set to zero ($w=0$) for the QVA algorithm are also included since, in the analysis shown in [22], for most large circuits (over 100 two-qubit gates), the optimal $w$ was close to 0.
Mean Absolute Mean Squared $R^{2}$ Score $\uparrow$ Pearson
Error (MAE) $\downarrow$ Error (MSE) $\downarrow$ Correlation $\uparrow$
Qiskit 0.422 0.273 -0.831 0.606
Qiskit +$T_{1,2}$ 0.094 0.024 0.843 0.927
ESP 0.229 0.098 0.341 0.827
ESP +$T_{1,2}$ 0.080 0.016 0.891 0.945
QVA 0.134 0.034 0.769 0.903
QVA +$T_{1,2}$ 0.161 0.051 0.658 0.898
QVAw=0 0.126 0.043 0.715 0.895
QVAw=0 +$T_{1,2}$ 0.088 0.015 0.899 0.951
This Work 0.210 0.084 0.434 0.849
This Work +$T_{1,2}$ 0.067 0.008 0.944 0.976
The results presented in Table 1 highlight the performance of the proposed model (including $T_{1}$ and $T_{2}$), which consistently outperforms all other approaches across the selected metrics. It achieves the lowest MAE, improving between 15.63% and 84.07% over other techniques; the lowest MSE, with improvements ranging from 44.14% to 96.92%; the highest $R^{2}$ score, with gains of 4.96% to 213.54%; and the highest Pearson correlation, improving by 2.59% to 61.05%. The results including $T_{1,2}$ are a significant improvement across almost all models, demonstrating the importance of incorporating coherence noise into the fidelity estimation models. This suggests that including additional noise factors can significantly enhance the model’s accuracy in predicting the fidelity of quantum circuits executed on a real device, making it a highly promising method for future quantum error modeling and prediction.
Since the QVA model already predicts a lower fidelity than the observed success rate even without accounting for $T_{1}$ and $T_{2}$ errors, incorporating coherence errors further worsens its predictions. This highlights the need for a stronger theoretical framework for more accurate fidelity estimation.
To ensure the robustness of our results, we repeated the experiments multiple times, conducting a total of six independent runs executed in different days. The proposed model consistently achieved the best performance across different calibration instances and quantum processors in all cases. Figure 8 provides a detailed comparison of the prediction differences observed across all estimation methods, as well as the obtained $R^{2}$ scores for each experiment.
@@ 5 Results
The underlying code for this study and the data generated and analyzed are available in the GitHub repository, https://github.com/escofetpau/Analytic-Model-of-Fidelity-under-Depolarizing-Noise.
6 Discussion
The performance differences between the considered fidelity estimation models reveal key insights into their underlying assumptions and applicability. The QVA model, despite showing competitive performance in regression metrics, predicts lower fidelities than those observed in real hardware for some circuits. This behaviour, combined with the exclusion of significant error sources such as coherence or crosstalk noise, suggests incomplete theoretical foundations. While the QVA model may provide reasonable predictions in some contexts, its tendency to deviate from hardware-observed outcomes raises questions about its reliability in diverse scenarios. It is, therefore, crucial not only to evaluate models based on final metrics but also to examine their assumptions and mechanisms, ensuring they align with the complexities of real devices.
Unlike simulations requiring extensive computational resources, the scalable depolarizing model proposed in this work can efficiently and accurately predict circuit fidelities, opening new avenues for exploration. For instance, it enables comparative studies of compilation strategies seeking to maximize computation fidelity, as showcased in existing literature [22]. Additionally, the model facilitates the exploration of computational bounds for varying architectural parameters. By adjusting single- and two-qubit depolarizing factors ($p_{1}$ and $p_{2}$), the model can determine the conditions necessary to achieve high-fidelity computation. Figure 6 serves as an example. It illustrates, for various quantum circuits, the required $p_{2}$ values for achieving a circuit fidelity above 0.99 as the number of qubits increases for fixed $p_{1}$ values.
These types of analysis are only possible thanks to the proposed estimation method’s light scalability and high accuracy. If the model were inaccurate, it would lead to misleading predictions, potentially guiding hardware design and algorithmic decisions in the wrong direction. Additionally, without it being scalable, it would be unfeasible to thoroughly explore the full range of architectural parameters or extend the analysis to a higher number of qubits. By enabling precise and efficient fidelity estimations, the proposed approach allows for an in-depth evaluation of different quantum computing scenarios, supporting the development of more reliable and optimized quantum hardware and software.
7 Conclusions
Fidelity serves as the ultimate and most reliable metric for assessing the performance of quantum computations, directly reflecting the impact of noise and imperfections on quantum circuits. Accurate fidelity estimation is crucial for benchmarking quantum processors, optimizing compilation strategies, and guiding architectural decisions.
In this work, we have shown that state-of-the-art fidelity estimation techniques often fail to predict the performance of real-world quantum processors accurately. The discrepancies observed between predictions and actual hardware executions highlight the limitations of existing approaches, which either oversimplify noise models or lack scalability for large-scale quantum systems.
To address these challenges, we developed a theoretical framework that analytically characterizes the effects of depolarizing noise on quantum circuit fidelity. The proposed model provides an efficient and scalable approach to fidelity estimation, leveraging device calibration data to offer a more precise assessment of quantum computation reliability.
We validated our method against alternative estimation techniques and real hardware executions on IBM Q processors, consistently achieving the most accurate predictions across all tested scenarios. These results confirm the robustness of our approach in practical quantum computing environments.
Beyond its immediate application in fidelity estimation, our model’s scalability enables a wide range of use cases, including hardware design optimization, performance evaluation of emerging quantum architectures, and feasibility analysis of large-scale computations. By providing a reliable and computationally efficient methodology, our work contributes to the advancement of quantum computing by offering a powerful tool for designing and assessing both quantum processors and quantum circuits.
Acknowledgements
@@ Appendix A Robustness of Fidelity Estimation Across Multiple Hardware Executions
Figure 8: Comparison of the fidelity prediction accuracy for five different estimation models, each incorporating coherence errors ($T_{1}$ and $T_{2}$). The figure consists of five rows, one per model (including QVA with $w=0$), and six columns, each corresponding to an independent experiment conducted on different days using IBM Q Eagle processors (IBM Q Kyiv, IBM Q Sherbrooke, and IBM Q Brisbane). Each cell represents a specific experiment’s fidelity error for each model. The $R^{2}$ metric is reported for each model, quantifying the correlation between predicted and measured fidelities. Across all experiments, the proposed model consistently achieves the highest $R^{2}$ values, ranging from $0.87$ to $0.94$, demonstrating superior predictive accuracy and robustness.
Appendix B Software Availability
To ensure full reproducibility of our results, all code, quantum circuits (in QASM format), generated data, and analysis scripts used in this study are publicly available in our GitHub repository: https://github.com/escofetpau/Analytic-Model-of-Fidelity-under-Depolarizing-Noise.
This repository contains the source code for all fidelity estimation methods, including Qiskit Simulation, ESP, QVA, our proposed approach, and all circuits used in our experiments. Additionally, it provides six sets of experimental data collected from real hardware executions and the necessary scripts to regenerate all figures in the paper. By making our methodology accessible, we aim to facilitate independent validation and further research on fidelity estimation and quantum hardware performance analysis.